OpenStudy (anonymous):

2x-y=4,find the value of m,b and the x-intercept

OpenStudy (anonymous):

Okay, slope-intercept form: y = mx + b 2x - y = 4 -y = -2x + 4 y = 2x - 4 m = 2 b = 4 m*b = 8 x intercept: 0 = 2x - 4 -2x = -4 x = 2 (2,0) is our x intercept.

OpenStudy (anonymous):

wow.! thank you so much for your help.

OpenStudy (anonymous):

No problem. :D